OpenStudy (anonymous):

what is the equation in point-slope form of the line passing through (0,6) and (1,3)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 0}}\quad ,&{\color{blue}{ 6}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\Leftarrow\textit{point-slope form, plug in your values}\)

OpenStudy (anonymous):

(y-3)=3(x+6)

OpenStudy (anonymous):

am i right?

OpenStudy (jdoe0001):

ahemm what did you get for the slope?

OpenStudy (anonymous):

-3 and 6

OpenStudy (jdoe0001):

hmm 6? well, -3 is right.. yes \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 0}}\quad ,&{\color{blue}{ 6}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ 6}}=\color{green}{ -3}(x-{\color{red}{ 0}})\)

OpenStudy (jdoe0001):

or y-6=-3x

OpenStudy (anonymous):

so i was right? @jdoe0001

OpenStudy (jdoe0001):

well... you had -> (y-3)=3(x+6) <- the slope is correct, is -3 the (y-3)=3(x+6) ^ ^ indicates a point at (-6, 3) which you don't have

OpenStudy (anonymous):

oh so it would be (y-3)=-3(x-1)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ \large \bf 0}}\quad ,&{\color{blue}{\large \bf 6}})\quad &({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ \large y-{\color{blue}{ 6}}=\color{green}{ -3}(x-{\color{red}{ 0}})\)

OpenStudy (jdoe0001):

you can really just use either... so if we use the 2nd point you'd end up with -> (y-3)=-3(x-1) <-- yes and that's valid too

OpenStudy (jdoe0001):

since both equations will simplify to the same

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

yw